# DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing

### v2

@@ 5.2. Game Review Analysis
Metric Baseline Optimized
Hallucination Rate (lower is better) 0.465 0.312
Sentiment Accuracy (higher is better) 0.664 0.650
Kendall’s Tau (higher is better) 0.470 0.631
@@ 5.2.1. Implementations
Since the documents exceed context limits and require temporal reasoning, we do not compare against existing systems, which do not support documents beyond context windows. Our baseline DocETL pipeline consists of a single map to extract positive_reviews and negative_reviews (both list types), with documents truncated from the middle to fit the context window—effectively randomly sampling reviews from each game’s corpus. The operation looks like the following:
@@
Given the following reviews for the game {{ input.app_name }}, analyze them and select 10 positive and 10 negative reviews that are evenly distributed across time:
{{ input.concatenated_reviews }}
@@
DocETL’s optimizer transforms this pipeline into: (a) A split operation that chunks input by token count (104,652 tokens per chunk): no gather operation (b) Two map operations per chunk—one each for positive/negative reviews—each incorporating one round of gleaning (directive 7) to ensure that the reviews are valid (c) A reduce operation to combine the positive and negative reviews from the chunks and present them in chronological order, matching the original get_reviews operation’s output schema.
@@ 5.2.2. Results
As in Table 4, the optimized pipeline significantly reduced hallucinations and improved temporal ordering. We observe a 32.9% reduction in hallucination rate (from 46.5% to 31.2%), demonstrating more reliable review extraction. Sentiment accuracy remained relatively stable (66.4% vs 65.0%), suggesting that improved extraction did not come at the cost of sentiment classification quality. We also have a 34.3% improvement in Kendall’s Tau (from 0.470 to 0.631), i.e., better temporal ordering.
The optimized pipeline costs \$1.48 (173.63s runtime) versus the baseline’s \$0.12 (29.27s). However, the baseline achieves this by truncating data to fit LLM context limits. With full data processing, the baseline would cost \$0.28, making the optimized pipeline 5.3×5.3$\times$5.3 × more expensive—still less than the 10×10$\times$10 × cost gap between gpt-4o and gpt-4o-mini models. This cost increase is justified by the improved temporal reasoning accuracy, and is due to steps like gleaning (which doubles operation cost); however, the gleaning validator consistently flagged temporal issues; in nearly all instances, the validator provided feedback like “The … reviews are not sorted correctly by timestamp; they should be organized chronologically.” And, as mentioned previously, costs are dropping precipitously. The optimization cost was \$6.60; however, this is a one-time cost.
@@ 5.3. Declassified Article Analysis
We evaluate DocETL’s effectiveness on resolve and reduce tasks using 733 paranormal case files from The Black Vault, a repository of declassified international government documents, averaging 700 words each. Each article documents a reported paranormal event with details such as location and witness accounts. We scraped articles from their website and used Azure Document Intelligence to convert any PDF attachments to text, and provide this data for transparency. [^5] https://osf.io/9xsbq Our task is to determine the distinct locations for each type of paranormal event (e.g., all cities and regions where UFO sightings were reported). The task involves two challenges: (i) standardizing event types across articles, and (ii) extracting and aggregating location mentions across articles for each event type.
We evaluated precision of extracted locations by first programatically verifying their presence in the source text and attempting to geocode them using the Nominatim API, based on OpenStreetMap. For locations in the text that could not be geocoded (e.g., specific rivers or mountain ranges), we performed manual verification.
@@ 5.3.1. Implementations
We consider 3 pipelines. We only consider one baseline, written in DocETL, as other systems don’t have a resolve operator. This pipeline consists of: (i) a map to extract the event type (e.g., “humanoid sighting”) per article, and (ii) a reduce to aggregate distinct locations across all articles of each event type.
DocETL’s optimizer modified this pipeline in two ways. First, it synthesized a resolve between map and reduce to standardize event types (directive 9). Second, it optimized reduce by determining an appropriate fold batch size (41) to process document batches, synthesizing the corresponding fold prompts. The optimized pipeline consists of: (i) a map (as before), (ii) a resolve to standardize event types (e.g., variations of “UFO sighting”), and (iii) a reduce (as before), but using a batched fold, with batch size 41.
To isolate the impact of the optimized reduce operation, we also evaluate the a version of this pipeline (+resolve only), which uses the original reduce operation without batched folding.
@@ 5.3.2. Results
Metric Baseline +Resolve Only Optimized
Location Precision N/A 0.994 1.000
Location Recall N/A 298 435
Distinct Event Types 164 83 83
As shown in Table 5, the baseline pipeline extracts 233 distinct event types with many semantic duplicates (e.g., “UFO Sighting”, “Category: UFO Sighting”, “Event Type: UFO Sighting”), making location aggregation impractical as most event types contain only one article. Adding resolve enables meaningful aggregation by consolidating to 83 event types. The +resolve only pipeline extracts 298 locations with 99.4% precision, and the optimized pipeline further improves this to 100% precision while extracting 435 locations (46% higher recall). This improved recall arises because batched folding allows the LLM to incrementally process and track distinct locations, rather than attempting to process all documents at once, where important details may be lost due to LLM context window overload (Levy et al., 2024; Liu et al., 2024a).
Cost Analysis. The resolve-only pipeline cost \$1.16 (307.36s runtime), while the optimized version cost \$1.34 (625.64s runtime), plus \$0.50 for optimization. The longer runtime for the optimized pipeline stems from using multiple LLM calls per event type during folding, compared to the resolve-only version which makes just one LLM call to process all documents per event type. For all operations, and the optimizer agents, we used gpt-4o-mini.
@@ 5.4. Biomedical Classification
We evaluate DocETL and LOTUS (Patel et al., 2024) on a challenging biomedical drug reaction classification task (D’Oosterlinck et al., 2023), introduced in the LOTUS preprint. For each of 250 biomedical papers, the task involves identifying which out of a canonical list, called the MedDRA list, of 24,300 adverse drug reactions, are discussed, amounting to a multi-label classification problem. Performance is measured using rank-precision@k (RP@k), evaluating both the accuracy of identified reactions and their ranking—a higher score indicates that true positive reactions appear earlier in the list.
@@ 5.4.1. Implementations
For LOTUS, we report numbers from their preprint. In addition, to evaluate against LOTUS on an equal footing using the same models (gpt-4o-mini for LLM calls, text-embedding-3-small for embeddings), maintaining similar LLM call budgets (12k calls), we modified their codebase to replicate the hand-written pipeline from their preprint to our best ability. We present these numbers as well.
@@
LOTUS Baseline. While LOTUS’s paper describes an efficient map-search-filter pattern to implement semantic joins, their codebase (as of November 2024) implements joins as a simple nested loop that invokes an LLM for every tuple pair with the user-defined natural language predicate; i.e., “The article {left:article} indicates that the patient is experiencing the {right:reaction} in our case. This would require over 6 million LLM calls on our dataset, which is prohibitively expensive. Therefore, we implemented their described map-search-filter pattern ourselves as a pandas dataframe accessor: first extracting reactions from each article via a map operation, then using similarity search to find candidate MedDRA labels, and finally filtering these candidates with an LLM to verify matches.
@@
Our implementation follows their described pipeline with some necessary deviations. First, a map operation extracts drug reactions from each article (one LLM call per article). Then, for each article’s extracted reactions, we find the 49 nearest neighbors in embedding space among the MedDRA labels—we chose k=49 to match our target LLM call budget ($\sim 12k$, equivalent to the number of LLM calls executed by DocETL). While LOTUS uses FAISS for similarity search (Douze et al., 2024), we encountered segmentation faults with their implementation. We instead use exact cosine similarity computation via NumPy, which may slightly increase runtime but provides more precise results—though this overhead is minimal given we only search through O(10k) vectors per article on an M1 Mac. Our filter operation has access to the full article text rather than just the map operation output, making our implementation generous to LOTUS.
@@
To evaluate rank-precision, we post-processed the pipeline outputs into ranked lists. For each article, we ranked the matching labels based on their semantic similarity scores (obtained during the similarity search phase), taking the top 5 or 10 labels for RP@5 and RP@10 computation respectively.
@@
DocETL Implementation. In DocETL, we implement this task as an equijoin between articles and MedDRA labels, using a comparison prompt that asks “Can the following condition be found in the article?” The prompt includes both the article text and the label, as well as an indicator of whether the condition text appears as a substring of the article (which is possible in Jinja templating). We do not evaluate an unoptimized version of this pipeline, due to the over 6 million LLM call estimate (24,300 labels $\times$ 250 articles). DocETL optimized this into a map-equijoin pipeline. The map extracts the medical conditions per article, with a prompt designed for medical text but without demonstrations or examples. For equijoin, DocETL synthesized blocking rules including an embedding similarity threshold of 0.5253 and a requirement that all words in the reaction label appear in the article text. Finally, we add a reduce operation that asks the LLM to rank the identified labels for each article from most to least confident, enabling evaluation of ranking quality. To maintain a fair comparison with LOTUS, we did not apply DocETL’s reduce operator optimizations to this ranking step.
@@ 5.4.2. Results
System RP@5 RP@10 RP@25
DocETL 0.281 0.313 0.371
LOTUS (Our reimplementation) 0.213 0.207 0.206
LOTUS (Reported) 0.241 0.258 N/A
The difference in ranking approaches does not disadvantage either system. At RP@25, which effectively measures recall since articles contain fewer than 25 relevant labels in the ground truth, DocETL still shows an 80% improvement over reimplemented LOTUS. For RP@5 and RP@10, DocETL shows 33% and 50% improvements, respectively. This substantial improvement in recall likely stems from DocETL’s synthesized blocking rules: while LOTUS relies purely on embedding similarity to identify candidate reactions, DocETL’s additional rule requiring all words in the reaction label to appear in the article text may surface relevant reactions that have low embedding similarity scores. While LOTUS’ reported performance is greater, it still falls well short of DocETL. We suspect this gap may be because the reported version provides multiple few-shot examples—which we don’t.
Cost and Dataset Analysis. Our reimplemented LOTUS pipeline costs \$0.47 and takes 925 seconds to run. The DocETL pipeline costs \$3.65 and takes 463.28 seconds, with an additional optimization cost of \$2.37. We believe this additional cost is an acceptable tradeoff because of the between 30-80% improvement in RP; moreover, these costs would be insignificant with open-source LLMs. The runtime can be highly variable; as LLMs can have high tail latencies, and LOTUS and DocETL may implement LLM retry logic differently.
@@ 5.5. Case Study: Police Misconduct
This pipeline begins with a map operation to extract metadata ($\text{Map}_{M}$), followed by document chunking of 12840 tokens each (Split). Each chunk then undergoes parallel processing: $\text{Map}_{S}$ for summarization and $\text{Map}_{H}$ for header extraction. The Gather operation collects context for each chunk, including the header lineage for the current chunk, 2 full previous chunks, and summaries of the other previous chunks. The original Map operation is then applied to each rendered chunk, with gleaning applied for refinement. Results from all chunks of a document are combined using $\text{Reduce}_{D}$. The pipeline then flattens the results (Unnest), resolves officer names (Resolve), and finally summarizes misconduct per officer (Reduce). This optimized pipeline incorporates several of our rewrite rules, including document chunking (1), header lineage context and summarization (3), gleaning for the Map operations (7), and duplicate key resolution (9).

### v3

@@ 5.2. Game Review Analysis
Metric DocETL (Unopt.) DocETL (Opt.) Non-LLM
Hallucination Rate (lower is better) 0.465 0.312 N/A
Sentiment Accuracy (higher is better) 0.664 0.650 0.605
Kendall’s Tau (higher is better) 0.470 0.631 N/A
@@ 5.2.1. Implementations
Since the documents exceed context limits and require temporal reasoning, we do not compare against existing LLM-based systems, which do not support documents beyond context windows. Our baseline DocETL pipeline consists of a single map to extract positive_reviews and negative_reviews (both list types), with documents truncated from the middle to fit the context window—effectively randomly sampling reviews from each game’s corpus. The operation looks like the following:
@@
Given the following reviews for the game {{ input.app_name }}, analyze them and select 10 positive and 10 negative reviews that are evenly distributed across time: {{ input.concatenated_reviews }}
@@
DocETL’s optimizer transforms this pipeline into: (a) A split operation that chunks input by token count (104,652 tokens per chunk): no gather operation (b) Two map operations per chunk—one each for positive/negative reviews—each incorporating one round of gleaning (directive 7) to ensure that the reviews are valid (c) A reduce operation to combine the positive and negative reviews from the chunks and present them in chronological order, matching the original get_reviews operation’s output schema. We added a non-LLM baseline that extracts reviews via regex, classifies sentiment with NLTK and VADER (Bird et al., 2009; Hutto and Gilbert, 2014), and selects the first 10 positive and negative reviews. Since this baseline only performs classification (i.e., it is not a generative model), hallucination rate and Kendall’s Tau metrics don’t apply.
@@ 5.2.2. Results
As shown in Table 4, we observe a 32.9% reduction in hallucination rate (from 46.5% to 31.2%), demonstrating more reliable review extraction. Sentiment accuracy remained stable (66.4% vs 65.0%), while Kendall’s Tau improved by 34.3%, indicating better temporal ordering. Both LLM-based approaches outperform the non-LLM baseline in sentiment accuracy, despite having to handle complex additional tasks beyond simple sentiment classification.
The optimized pipeline costs \$1.48 (173.63s runtime) versus the baseline’s \$0.12 (29.27s). However, the baseline achieves this by truncating data to fit LLM context limits. With full data processing, the baseline would cost \$0.28, making the optimized pipeline 5.3×5.3$\times$5.3 × more expensive—still less than the 10×10$\times$10 × cost gap between gpt-4o and gpt-4o-mini models. This cost increase is justified by the improved temporal reasoning accuracy, and is due to steps like gleaning (which doubles operation cost); however, the gleaning validator consistently flagged temporal issues; with feedback like “The … reviews are not sorted correctly by timestamp; they should be organized chronologically.” The optimization cost was \$6.60; however, this is a one-time cost. The non-LLM baseline had a runtime of 15.89 seconds.
@@ 5.3. Declassified Article Analysis
We evaluate DocETL’s effectiveness on resolve and reduce tasks using 733 paranormal case files from The Black Vault, a repository of declassified international government documents, averaging 700 words each. Each article documents a reported paranormal event with details such as location and witness accounts. We scraped articles from their website and used Azure Document Intelligence to convert all PDF attachments to text, and provide this data for transparency at https://osf.io/9xsbq. Our task is to determine the distinct locations for each type of paranormal event (e.g., all cities and regions where UFO sightings were reported). The task involves two challenges: (i) standardizing event types across articles, and (ii) extracting and aggregating location mentions across articles for each event type.
We evaluated precision of extracted locations by first programatically verifying their presence in the source text and attempting to geocode them using the Nominatim API, based on OpenStreetMap. We also measured hallucination rate—a subset of precision—defined as the proportion of locations that don’t exist in the source text. For locations in the text that could not be geocoded (e.g., specific rivers or mountain ranges), we performed manual verification.
@@ 5.3.1. Implementations
We consider 4 pipelines. We only consider one LLM-powered baseline, written in DocETL, as other systems don’t support resolve. This pipeline consists of: (i) a map to extract event type (e.g., “humanoid sighting”) per article, and (ii) a reduce to collect distinct locations across all articles of each event type. DocETL’s optimizer modified this pipeline in two ways. First, it synthesized a resolve between map and reduce to standardize event types (directive 9). Second, it optimized reduce by determining a fold batch size (41) to process document batches, synthesizing the corresponding fold prompts. The optimized pipeline consists of: (i) a map (as before), (ii) a resolve to standardize event types (e.g., variations of “UFO sighting”), and (iii) a reduce (as before), but using a batched fold, with batch size 41. To isolate the impact of the optimized reduce operation, we also evaluate the a version of this pipeline (+resolve only), which uses the original reduce operation without batched folding. Our 4th pipeline represents a non-LLM baseline that extracts location (LOC) entities from article text using spaCy’s en_core_web_lg model (Honnibal et al., 2020). This script processes the resolved results from DocETL’s optimized pipeline to establish a comparison point for location precision and recall.
@@ 5.3.2. Results
Metric DocETL (Unopt.) DocETL (+Resolve Only) DocETL (Opt.) Non-LLM
Location Precision N/A 0.994 1.000 0.6812
Location Recall N/A 298 435 435
Distinct Event Types 164 83 83 N/A
Hallucination Rate N/A 0.01 0.01 0.00
As shown in Table 5, the baseline DocETL pipeline extracts 233 distinct event types with many semantic duplicates (e.g., “UFO Sighting”, “Category: UFO Sighting”, “Event Type: UFO Sighting”), making location aggregation impractical as most event types contain only one article. Adding resolve enables meaningful aggregation by consolidating to 83 event types. The +resolve only pipeline extracts 298 locations with 99.4% precision, and the optimized pipeline further improves this to 100% precision while extracting 435 locations (46% higher recall). The non-LLM baseline matches the optimized pipeline’s recall but with substantially lower precision (68.12% vs. 100%), highlighting the LLM’s superior ability to accurately identify relevant locations in context. All systems exhibit low hallucination rates. This improved recall arises because batched folding allows the LLM to incrementally process and track distinct locations, rather than attempting to process all documents at once, where important details may be lost due to LLM context window overload (Levy et al., 2024; Liu et al., 2024a).
Cost Analysis. The resolve-only pipeline cost \$1.16 (307.36s), while the optimized version cost \$1.84 (\$1.34 + \$0.50 for optimization; 625.64s). The optimized pipeline’s longer runtime results from multiple LLM calls per event type during folding, versus one call in the resolve-only version. The non-LLM baseline ran in 158.85s. For all operations, and the optimizer agents, we used gpt-4o-mini.
@@ 5.4. Biomedical Classification
We evaluate DocETL on the challenging Biodex biomedical drug reaction classification task from the LOTUS paper (Patel et al., 2024). For each of 250 biomedical papers, the task involves identifying which of 24,300 adverse drug reactions (from the MedDRA list) are discussed. Performance is measured using rank-precision@k (RP@k), evaluating both accuracy and ranking of identified reactions. A higher score indicates that true positive reactions appear earlier in the list. We also evaluate the hallucination rate, measuring the proportion of identified reactions that are not present in the drug reaction list.
@@ 5.4.1. Implementations
We compare against LOTUS using numbers from their preprint and our reimplementation of their pipeline using the same models (gpt-4o-mini for LLM calls, text-embedding-3-small for embeddings) for fair comparison.
@@
LOTUS Baseline. We implemented their map-search-filter pattern as a pandas dataframe accessor: first extracting reactions from each article via a map operation, then using similarity search to find candidate MedDRA labels, and finally filtering these candidates with an LLM to verify matches.
@@
Our implementation follows their described pipeline with some necessary deviations. First, a map operation extracts drug reactions from each article (one LLM call per article). Then, for each article’s extracted reactions, we find the 49 nearest neighbors in embedding space among the MedDRA labels—we chose k=49 to match our target LLM call budget ($\sim 12k$, equivalent to the number of LLM calls executed by DocETL). We use exact cosine similarity computation via NumPy instead of a FAISS approximate nearest neighbor index, which may slightly increase runtime but provides more precise results—though this overhead is minimal given we only search through O(10k) vectors per article on an M1 Mac.
@@
To evaluate rank-precision, we post-processed the pipeline outputs into ranked lists. For each article, we ranked the matching labels based on their cosine similarity scores (obtained during the similarity search phase), taking the top k labels for RP@k computation.
@@
DocETL Implementation. In DocETL, we implement this task as an equijoin between articles and MedDRA labels, using a comparison prompt that asks “Can the following condition be found in the article?” The prompt includes both the article text and the label, as well as an indicator of whether the condition text appears as a substring of the article (which is possible in Jinja templating). We don’t evaluate an unoptimized version due to the impractical number of LLM calls required (over 6 million). DocETL optimized this into a map-equijoin pipeline, where the map extracts medical conditions per article(with a prompt designed for medical text but without demonstrations or examples), and the equijoin uses synthesized blocking rules including an embedding similarity threshold of 0.5253 and a requirement that all words in the reaction label appear in the article text. Finally, we add a reduce operation that asks the LLM to rank the identified labels for each article from most to least confident, enabling evaluation of ranking quality. We did not apply DocETL’s reduce operator optimizations to this ranking step.
We also include a simple non-LLM baseline that identifies candidate labels by checking for exact substrings and ranks them by length. Given the large number of required comparisons, we opted for a keyword baseline over more complex NLP libraries (e.g., NTLK).
@@ 5.4.2. Results
System RP@5 RP@10 RP@25 Hallucination
DocETL 0.281 0.313 0.371 0.001
LOTUS (Our reimplementation of map-search-filter) 0.213 0.207 0.206 0.000
LOTUS (Reported) 0.241 0.258 N/A 0.000
Non-LLM Baseline 0.106 0.158 0.262 0.000
At RP@25, which effectively measures recall since articles contain fewer than 25 labels in the ground truth, DocETL shows an 80% improvement over reimplemented LOTUS. For RP@5 and RP@10, DocETL shows 33% and 50% improvements, respectively. The non-LLM baseline achieves lower RP@5 and RP@10 scores than the LLM-based methods, but a competitive RP@25 (still worse than the plan DocETL found). The improvement in recall likely stems from DocETL’s synthesized blocking rules: while LOTUS relies purely on embedding similarity to identify candidate reactions, DocETL’s rule requiring all words in the reaction label to appear in the article text may surface reactions that have low embedding similarity scores. In terms of hallucination rate, all systems perform well with essentially zero hallucinations, though DocETL has a marginally higher rate of 0.001 on average. The difference between LOTUS’ reported performance and our reimplementation may be attributed to model choices and prompting strategies, as we standardized on gpt-4o-mini without “few-shot” examples for consistency across all experiments.
Cost and Dataset Analysis. The non-LLM baseline takes 290.65 seconds to run. Our reimplemented LOTUS pipeline costs \$0.47 and takes 925 seconds to run. The DocETL pipeline costs \$3.65 and takes 463.28 seconds, with an additional optimization cost of \$2.37. We believe this additional cost is an acceptable tradeoff because of the between 30-80% improvement in RP; moreover, these costs would be insignificant with open-source LLMs. The runtime can be highly variable; as LLMs can have high tail latencies, and LOTUS and DocETL may implement LLM retry logic differently.
@@
Figure 5. Cost vs. metrics (precision, recall, and F1) for 30 different LLM-generated implementations of rewrite directives applied to the legal contract analysis task. Each point represents a distinct plan implementation, colored by directive type; isolated projections (Equation 11, chaining projections (Equation 10, or gleaning (Equation 7). The DocETL unoptimized baseline and optimized plan from Section 5.1 are shown with dashed lines for reference, though not generated in this experiment. Due to the optimizer’s nondeterministic nature, some plans in this experiment achieved higher metrics than the original optimized plan.
@@ 5.5. Case Study: Police Misconduct
This pipeline begins with a map operation to extract metadata ($\text{Map}_{M}$), followed by document chunking of 12840 tokens each (Split). Each chunk then undergoes Directive Equation 3: $\text{Map}_{S}$ for summarization and $\text{Map}_{H}$ for header extraction. The Gather operation collects context for each chunk, including the header lineage for the current chunk, 2 full previous chunks, and summaries of the other previous chunks. The original Map operation is then applied to each rendered chunk, with gleaning applied for refinement. Results from all chunks of a document are combined using $\text{Reduce}_{D}$. The pipeline then flattens the results (Unnest), resolves officer names (Resolve), and finally summarizes misconduct per officer (Reduce). Overall, this optimized pipeline incorporates several of our rewrite rules, including document chunking (1), header lineage context and summarization (3), gleaning for the Map operations (7), and duplicate key resolution (9).
@@
5.6. Case Study: Rewrite Directives in Legal Contract Analysis
One may wonder: if DocETL’s agentic optimizer depends on how effectively LLMs transform abstract rewrite directives into concrete plans, how well do they actually perform? We examined this with a case study on our legal contract analysis task from Section 5.1. We selected three rewrite directive types that don’t require physical parameter tuning: projection chaining (synthesizing a chain of dependent map operations, according to Equation 10), isolated projection (synthesizing independent map operations, per Equation 11), and gleaning (Equation 7). Using our GPT-4o-powered LLM agent, we generated 10 different instantiations of each directive, yielding 30 distinct plans with varying prompts, structures, and decomposition strategies.
Our study had two objectives: (i) to assess the quality of LLM-implemented directives and (ii) to evaluate our LLM-based plan evaluation mechanism. We executed all 30 plans on the 50 contracts, measuring precision, recall, F1 scores, and cost to characterize performance distribution across implementations of the same directive. We then had our optimizer (using GPT-4o-mini as judge) rank these plans per the approach in Section 4, and computed the Kendall’s Tau between these rankings and actual performance metrics.
Figure 5 illustrates the results. We start with (i). Projection synthesis rewrites (both chaining and isolation) showed high variance in both cost and accuracy metrics. Cost variance was expected: more projections synthesized means more LLM calls and proportionally higher costs. More surprising was the substantial accuracy variance, which stemmed from implementation inconsistencies. 20% of the generated plans (all from chain projection implementations) failed to include document placeholders in prompts (e.g., writing extract X clause from the document instead of extract X clause from {{ input.document }}), resulting in the LLM receiving no text to analyze. Isolated projections consistently outperformed chaining projections for this task, likely because clause extraction subtasks are mostly independent. Gleaning directive implementations showed more consistent cost profiles, which is expected since they involve exactly three LLM calls per document (initial map, validation, and refinement). Despite implementation variance, many LLM-generated plans outperformed the unoptimized DocETL baseline, with 47% of plans achieving better precision, 67% of them achieving better recall, and 40% of them achieving better F1 scores. When compared against our optimized plan from the main experiment, 30% still achieved better precision, 10% achieved better F1 scores, though none improved on recall. Moving on to (ii), the Kendall’s tau correlation between LLM judge rankings and actual F1 score rankings was 0.642. So, while there is considerable variance in implementation quality, our selection mechanism can identify the more effective plans.
Overall, while LLMs can translate abstract directives into working plans, quality varies considerably, and naive implementations may contain critical errors. This underscores the importance of generating many plans and our LLM-based plan evaluation mechanism and suggests potential improvements through better prompt engineering or selecting models particularly adept at prompt rewriting.